Study the following functions

1)f(x) = Inx/x

2) f(x) = ex-x +3


1
Expert's answer
2021-09-21T17:42:47-0400

(1) Graph of "f(x) = \frac{ln(x)}{x}"



Here, domain of "f(x)" is:

"x>0,x\neq0 \\
\therefore x\in R^+"

Range of "f(x)" is:

"f(x)\in(-\infty,\frac{1}{e}]"


(2) Graph of "f(x) = e^x-x +3"





Here, domain of "f(x)" is:

"x\in R"

Range of "f(x)" is:

"f(x)\in[4,\infty)"

System